On the occasion of Deepawali festival each student of a class sends greeting cards to the others. If there are 20 students in the class then the total number of greeting cards exchanged by the students is?

Grade:12

2 Answers

sayanti
6 years ago
The required formula is 2 * nC2=n*(n-1) since number of freinds is 20. Then n=2 Required number of greetings card 20*(20-1)=20*19=380
sayanti
6 years ago
The required formula is 2 * nC2=n*(n-1) since number of freinds is 20. Then n=20 Required number of greetings card 20*(20-1)=20*19=380